OpenStudy (jjuden):

A flag pole casts a shadow that is 30 feet in length. A rope that is 40 feet in length is tied to the top of the flag pole and is extended to the edge of the shadow. Find the height of the flag pole to the nearest tenth.

OpenStudy (anonymous):

You know what the Pythagorean theorem is?

OpenStudy (chris911):

so 50 I guess

OpenStudy (anonymous):

Basically it is a^2 + b^@=c^2 30*30=900 40*40=1600 2500 √2500=50